OpenStudy (jacobbirddog):

Raisins sell for $3.50 per pound, and granola sells for $5.90 per pound. Terri bought some raisins and some granola. The total weight was 2.1 pounds and cost $8.79. Let p represent the number of pounds of raisins. Which equation represents the situation described? A. –2.4p + 12.39 = 8.79 B. –2.4p + 8.79 = 12.39 C. 3.5p + 5.9 = 8.79 D. 3.5p + 8.79 = 5.9

OpenStudy (anonymous):

What is important in the situation?

OpenStudy (jacobbirddog):

i have to find what equation rep's the situation that is discribed above

OpenStudy (anonymous):

That's the problem but if I were to solve this myself I would say "ok they state rasins are represented by p that's probably important" What number(s) are connected to p?

OpenStudy (jacobbirddog):

Raisins sell for $3.50 per pound, and granola sells for $5.90 per pound. Terri bought some raisins and some granola. The total weight was 2.1 pounds and cost $8.79. Let p represent the number of pounds of raisins. Which equation represents the situation described? A. –2.4p + 12.39 = 8.79 B. –2.4p + 8.79 = 12.39 C. 3.5p + 5.9 = 8.79 D. 3.5p + 8.79 = 5.9

OpenStudy (jacobbirddog):

i reposted the thing

OpenStudy (anonymous):

I know, I was asking you to tell me what you think goes in front of p. The important numbers in this are 3.5, 5.9, and 8.79

OpenStudy (jacobbirddog):

hmm

OpenStudy (jacobbirddog):

3.5

OpenStudy (anonymous):

yes, now what that we have determined that what are we looking for? what are we adding to this?

OpenStudy (jacobbirddog):

5.9

OpenStudy (anonymous):

Why?

OpenStudy (jacobbirddog):

?

OpenStudy (jacobbirddog):

sorry im very new to algebra and i don't get it :P

OpenStudy (anonymous):

I want to make sure you understand what we are discussing

OpenStudy (anonymous):

We are adding it to 3.5p because we want to get the total.

OpenStudy (anonymous):

So what would our equation be?

OpenStudy (jacobbirddog):

well our total is 8.79 so we just have to add ty so much!

OpenStudy (anonymous):

:)

OpenStudy (jacobbirddog):

crap none of the answers = 8.79 :P

OpenStudy (jacobbirddog):

what would it be?!?!

OpenStudy (jacobbirddog):

ugh can't solve it...

OpenStudy (anonymous):

c?